SOHO DAILY REPORT Operations Day - 343/2003 (09 December 03) Mission Day 2930 SOHO FLIGHT OPERATIONS TEAM, NASA/GSFC Ref. #: SH-FOT-DR-031209 ==================================================================== PASS #1 Madrid DSS54 0640-1205 PASS #2 Goldstone DSS16 1329-2250 ********************************************************************* PASS #1 Madrid DSS54 0640-1205 Anomalies: None NRT: None Activities Summary: - SSR Dump - Collected MDI-M (0846-0903) - Collected MDI-H (0903-1145) - OBT Distribution to SWAN, CELIAS - Collected Ranging Data - SSR SEF Reset (x2) - SSR Record - Executed OCD #1873: APME Z-axis Improvement Operations ------------------ PASS #2 Goldstone DSS16 1329-2250 Anomalies: G03-0800 1340 Commanding activities delayed 6 minutes on D16 due to problems with the exciter's sweep execution. DR #G103815. G03-0801 1449 Lost 4 minutes 36 seconds of unrecoverable TLM due to the TCP halting at D16. DR #G103816. G03-0802 1938 Received intermittent degraded TLM until 2250 due to RFI at D16. DR #G103819. NRT: 1443-1444 1456-2050 Activities Summary: - SSR Dump - Collected MDI-M (1446-2233) - New Control Star Slot 5/ Magnitude 6.55 - New Tracking Star Slot 4/ Magnitude 5.50 - Reaction Wheel Speed Monitoring Daily Adjustment - CRS Drift Estimation Update - RSL Table Upload - Collected Ranging Data - SSR SEF Reset (x2) - Executed CS-14: Loss of TM in High Rate ------------------ -- End of Report --